# Energy Efficient Transmission Parameters Selection Method Using Reinforcement Learning in Distributed LoRa Networks

### v1

@@ IV-C Experimental Results
IV-C1 Selection ratio of each transmission value
@@
As shown in Fig. 3, our proposed method can achieve the highest proportion of selecting the minimum power. In the case of $\epsilon$-greedy, it took more time to explore compared to the proposed method, resulting in an increased probability of selecting a larger transmission power. Moreover, the ADR-Lite had the lowest proportion of selecting the minimum power. This is due to the low accuracy of exploring optimal parameters and the tendency to select larger transmission power when a transmission fails once. Therefore, our proposed method is considered to have the best balance between exploration and exploitation among the compared methods.
@@ IV-C3 Energy Efficiency
V Conclusion
This paper proposed an autonomous distributed transmission parameter selection method using reinforcement learning to improve the energy efficiency in LoRa networks. We implemented the proposed method in a constructed real LoRa network and evaluated the energy efficiency and successful transmission rate performance. As a result, our proposed method showed the best performance in both metrics. Furthermore, the reinforcement learning-based methods performed better in both metrics compared to fixed allocation and ADR-Lite. This is considered because each LoRa ED could select smaller transmission power while avoiding channel congestion through learning. In our further work, we will explore more communication parameter selection, the fairness of each LoRa ED, and the detailed energy consumption for executing different algorithms.

### v2

@@ IV-C Experimental Results
IV-C1 Selection ratio of each TP level
@@
As shown in Fig. 3, our proposed method can achieve the highest proportion of selecting the minimum power. In the case of $\epsilon$-greedy, it took more time to explore compared to the proposed method, resulting in an increased probability of selecting a larger transmission power. Moreover, the ADR-Lite had the lowest proportion of selecting the minimum power. This is due to the tendency to select larger transmission power when a transmission fails once.
@@ IV-C3 Energy Efficiency
V Conclusion and Future Work
This paper proposed an autonomous distributed transmission parameter selection method using reinforcement learning to improve the energy efficiency in LoRa networks. We implemented the proposed method in a constructed real high-density LoRa network and evaluated the energy efficiency and successful transmission rate performance. As a result, our proposed method showed the best performance in both metrics. Furthermore, the reinforcement learning-based methods performed better in both metrics compared to fixed allocation and ADR-Lite. This is considered because each LoRa ED could select smaller transmission power while avoiding channel congestion through learning.
Future challenges include expanding the type of transmission parameters to be optimized, such as incorporating the selection of Spreading Factor (SF) and Bandwidth (BW), to further enhance energy efficiency. Additionally, increasing the number of selectable parameters or considering parameter combinations is expected to expand the search space, potentially slowing convergence. Therefore, improving the algorithm used in the initial exploration phase to efficiently handle a large search space is necessary. Furthermore, when implementing the proposed method on resource-constrained IoT devices, it is essential to conduct a more detailed analysis of the computational overhead and memory requirements to ensure feasibility.